OpenStudy (anonymous):

Help me find the equation of a circle with center (2,2); & passes thru the origin ?

OpenStudy (mertsj):

Find the distance from (2,2) to (0,0). That will be the radius of the circle. Use the distance formula.

OpenStudy (ybarrap):

$$ r^2 = (x-h)^2 + (y-k)^2\qquad(\star)\\ $$ Where <h,k> is the center of the circle. Now, you want the circle to also pass through the circle, so you want r to be the distance from the center of the circle to the origin: $$ r = \sqrt{2^2 + 2^2}=\sqrt{2\times 4}=2\sqrt{2} $$ So now you know, r and you also know h and k. Plug these into the equation above (indicated by a \(\star\) above) to find the equation of a circle in terms of the variables x and y.

OpenStudy (anonymous):

if it intersects axis at -1 and 3 does that mean I use the distance formula again ?

OpenStudy (mertsj):

We don't care where it intersects the x axis. All we want is the equation of the circle. We know the center is (2,2) and the radius is 2sqrt2 so radius squared is 8 and the equation is: (x-2)^2+(x-2)^2=8 and that is the answer to the problem which says to find the equation of a circle.

OpenStudy (anonymous):

No this is a new problem center is 1,2 and intersects x-axis at -1 and 3, does that mean I use the distance formula again using the -1 and 3 ?

OpenStudy (ybarrap):

Yes, that's what it means.

OpenStudy (mertsj):

That means you need to find the radius by finding the distance from (1,2) to (3,0)

OpenStudy (anonymous):

Oh okay, thank bro.

OpenStudy (mertsj):

yw